OpenStudy (imrankhan):

If 20+15=24 and 64+13=42, than 11+28=?

OpenStudy (anonymous):

imrankhan from which book you take this question? do you like to tell me?

OpenStudy (anonymous):

11+28=9

OpenStudy (anonymous):

its a terminological question@imrankhan...

OpenStudy (imrankhan):

If 20+15=24 and 64+13=42, than 11+28=? (A) 31 (B) 34 (C) 33 (D) 36 Seen on a online paper and answer is (D)36 but i don't know the mathod... :-(

OpenStudy (anonymous):

@TuringTest @terenzreignz @dan815